OpenStudy (anonymous):

Find the first 3 terms and the 20th term of the sequence that has the given n th term. an = (−1)n (n − 8)2

OpenStudy (anonymous):

Looks like this? \[a_n=(-1)^n(n-8)^2\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

?

jhonyy9 (jhonyy9):

a_1 =(-1)^1(1-8)^2 =(-1)(-7)^2 = -49 a_2 =(-1)^2(2-8)^2 =(+1)(-6)^2 =36 a_3 =(-1)^3(3-8)^2 =(-1)(-5)^2 = -25 a_20 =(-1)^20(20-8)^2 =(+1)(12)^2 = 144